# Restoring Missing Modes of 21cm Intensity Mapping with Deep Learning: Impact on BAO Reconstruction

### v1

@@ Abstract
In 21cm intensity mapping of the large-scale structure (LSS), regions in Fourier space could be compromised by foreground contamination. In interferometric observations, this contamination, known as the foreground wedge, is exacerbated by the chromatic response of antennas, leading to substantial data loss. Meanwhile, the baryonic acoustic oscillation (BAO) reconstruction, which operates in configuration space to "linearize" the BAO signature, offers improved constraints on the sound horizon scale. However, missing modes within these contaminated regions can negatively impact the BAO reconstruction algorithm. To address this challenge, we employ the deep learning model U-Net to recover the lost modes before applying the BAO reconstruction algorithm. Despite hardware limitations, such as GPU memory, our results demonstrate that the AI-restored 21cm temperature map achieves a high correlation with the original signal, with a correlation ratio of approximately $0.9$ at $k\sim 1{\rm h/Mpc}$. Furthermore, subsequent BAO reconstruction indicates that the AI restoration has minimal impact on the performance of the ‘linearized’ BAO signal, proving the effectiveness of the machine learning approach to mitigate the impact of foreground contamination. Interestingly, we demonstrate that the AI model trained on coarser fields can be effectively applied to finer fields, achieving even higher correlation. This success is likely attributable to the scale-invariance properties of non-linear mode coupling in large-scale structure and the hierarchical structure of the U-Net architecture.
@@ 1 Introduction
These Missing information can significantly impact cosmological measurements. For one, the loss of Fourier modes reduces the statistical constraining power. For baryonic acoustic oscillations (BAO), this loss also affects data post-processing. Most current BAO measurements in galaxy surveys utilize a technique known as BAO reconstruction [35, 36, 37, 38, 39, 40], which aims to mitigate the non-linear degradation of the BAO signal. Despite the technical differences among various reconstruction algorithms, these methods generally seek to solve the continuity equation to ‘linearize’ the BAO signature. Since this process occurs in the configuration space, the absence of substantial Fourier modes can severely influence the outcome. However, it is still possible to address this challenge by recovering information through the nonlinear coupling between different modes. For example, the tidal reconstruction method [41] develops a quadratic estimator from the tidal field to extract large-scale modes from small-scale perturbations. Yet, as a perturbative approach, its effectiveness is primarily limited to large scales.
With the widespread application of artificial intelligence (AI), Machine learning (ML) has become a powerful tool for processing large-scale structure data and analyzing cosmic dataset [42, 43, 44, 45, 46, 47, 48, 49, 50]. It has also proven to be highly effective in addressing high-dimensional astronomical problems, such as reconstructing BAO signals from dark matter density fields [51], isolating the EoR signals with complex frequency-dependent beam effects [52], and identifying ionization regions in images with instrument noise [53, 54]. Therefore, in this study, we aims to apply the deep learning technique, particularly the U-Net model, to recover the large-scale structure modes in Fourier regions affected by foreground contamination. Additionally, we will examine the impact of the AI-restored data on BAO reconstruction. The structure of this paper is as follows: Section 2 provides a brief overview of foreground contamination and its impact on BAO reconstruction. In section 3, we introduce our main methodology, including details of the simulation data for 21cm intensity mapping, the U-Net architecture, and the BAO reconstruction process. The results are presented in Section 4, and we conclude with a discussion in Section 5.
@@ 2 Foreground Contamination and the Impact on BAO Reconstruction
2.1 Foreground Wedge
In both single-dish and interferometric observations, the spectral smoothness of the foreground indicates that the LSS signal may be compromised within the Fourier region where the wavenumbers along the line-of-sight direction are small, i.e. $k_{\parallel}<k_{\rm int}$ where $k_{\parallel}$ represent the line-of-sight component of the wave vector $\mathbf{k}$. This commonly referred to as the intrinsic foreground. The exact value of $k_{\rm int}$ depends on the specific foreground removal techniques employed and specific instrumental systematics; for our analysis, we have adopted a reasonable value of $k_{\rm int}=0.1~{}{\rm h/Mpc}$. Furthermore, the chromatic response of interferometer complicates the situation [55, 56, 57] by introducing additional coupling between spectral and spatial information. This coupling can lead to foreground leakage from low $k_{\parallel}$ to high $k_{\perp}$, with the effect being particularly severe at high $k_{\perp}$, where $k_{\perp}$ is the transverse components of $k$.
@@ 2.1 Foreground Wedge
in the two-dimensional $k_{\parallel}-k_{\perp}$ plane. In this paper, we consider $\theta_{\rm FOV}=2^{\circ}$, at the redshift of $z=1$, and cosmological parameters as follows $\Omega_{m}=0.31,\Omega_{b}=0.048,\Omega_{\Lambda}=0.69,\sigma_{8}=0.82,h=0.68$. The combination of intrinsic foreground and the mode mixing wedge results in significant information loss, which, as will be discussed, substantially impacts BAO reconstruction.
Figure 2: Visual comparison between the true brightness temperature (upper panels), the observed temperature with intrinsic and foreground wedge removed (middle panels), and the U-Net restored temperature (lower panels). Different columns demonstrate maps at various resolutions, ranging from $64^{3}$ (first column) to $128^{3}$ (second), $256^{3}$ (third), and $320^{3}$ (last column). As detailed in the text, the AI model is trained at a resolution of $128^{3}$ and subsequently applied to maps of varying resolutions. All brightness temperatures are in unit of ${\rm mK}$.
2.2 BAO Reconstruction
Despite its robustness and enormous success in constraining cosmological parameters in LSS surveys, the baryonic acoustic oscillation is still subject to various systematics. One of the most well-understood effects is the broadening of the BAO peak caused by the non-linear evolution of the bulk velocity of LSS [58]. In the Fourier space, this manifests as a dampening of the BAO wiggles, which can be modeled by a simple exponential function. Compared to the linear BAO peak, this effect reduces the measurement accuracy of given LSS survey in extracting the sound horizon scale. Moreover, the same non-linear process also leads to the emergence of non-Gaussianity, which results in additional information leakage into higher order statistics and further decreases the overall constraining power of the measurements. Various techniques have been developed to reverse such process. For example, [59] proposed a reconstruction algorithm based on the linear Zel’dovich approximation which can partially reverse nonlinear degradation of BAO signal. This technique has been widely used in recent galaxy surveys [60, 61, 62, 63].
@@ 2.2 BAO Reconstruction
where $\mathbf{q}$ and $\mathbf{x}$ are Lagrangian and Eulerian coordinates of particles respectively, $\delta^{K}_{ij}$ is the Kronecker delta, the displacement potential and vector are defined with $\mathbf{\Psi}(\mathbf{q})=\mathbf{x}-\mathbf{q}=\nabla\phi(\mathbf{q})$. By solving equation (2.4), the reconstruction algorithm reverse such non-linear transformation and produces a reconstructed $\phi_{\rm rec}$ field on a grid that is close to the Lagrangian coordinates. In the following, we will denote the Laplacian of the reconstructed field $\delta_{\rm rec}=\nabla^{2}\phi_{\rm rec}$. Consequently, the BAO signature of the reconstructed field closely resemble that of the linear theory. Since these algorithms solve for the displacement field in the configuration space, in 21cm intensity mapping experiments, the missing modes due to the foreground contamination will inevitably impact the performance of the reconstruction.
@@ 3 METHOD
At low redshifts, neutral hydrogen predominantly resides within galaxies. To simulate mock observations for the 21cm intensity mapping, we begin with a fast N-body simulation to obtain the dark matter distribution and identify halos. Using the prescription outlined by [16], we then generate the HI mass distribution $M_{\rm HI}(\mathbf{x})$. Finally, we convert this distribution into the brightness temperature of the 21cm signal.
@@ 3.1 Data Preparation
We utilize the publicly available COLA (COmoving Lagrangian Acceleration) simulation code  [^1] https://bitbucket.org/tassev/colacode/src/hg/ [64], a hybrid approach integrating second-order Lagrangian Perturbation Theory (2LPT) and N-body algorithms, to rapidly simulate multiple realizations of dark matter distributions. Specifically, we conducted thirty realizations using the same set of cosmological parameters described previously but with varying initial seeds. Each simulation contains $1024^{3}$ particles within a cubic box measuring $800{\rm h^{-1}Mpc}$. The initial conditions for these simulations were set at a redshift of $z_{\rm init}=29$.
Using the dark matter distributions described above, we then generate the corresponding distributions of neutral hydrogen for each realization. Given the relatively low mass resolution of our dark matter simulations, the resolved halo masses are significantly larger than the minimum halo mass that can physically host HI. Consequently, our HI map consists of two distinct components: resolved halos and a subgrid contribution. For resolved halos, we employ the Robust Overdensity Calculation using K-Space Topologically Adaptive Refinement (ROCKSTAR) halo finder [65] to identify halo catalogs. This includes determining their locations, masses, and peculiar velocities. Furthermore, we account for the redshift space distortions (RSD) by adjusting the real space coordinates of halos according to their peculiar velocities $v_{r}=\textbf{v}\cdot\hat{\textbf{r}}$, along the line-of-sight (LoS) direction, which we designate as the $z$ axis. The redshift space position s is then defined by
@@
For subgrid contribution, we assume these low-mass halos trace the dark matter field. Therefore, with the position and the peculiar velocity of DM particles, we construct the redshift-space matter distribution using Eq (3.1). In each cell with volume $V_{\rm cell}$ and mass $M_{\rm cell}$ at redshift $z$, we assume the number of halos with mass $m$ follows a Poisson distribution, with the average number density determined by the halo mass function. For halos that are viralized at redshift $z^{\prime}$, the average number density can be calculated by [66]
@@
$\nu_{10}=\frac{\left[\delta_{sc}(z^{\prime})-\delta_{0}(\delta,z)\right]}{\sigma^{2}(m)-\sigma^{2}(M_{\mathrm{cell}})}.$ (3.3)
Here $\rho_{\mathrm{crit}}$ is the critical density, $\sigma(m)$ represents the variance of density fluctuations on mass $m$ and $\delta_{\text{sc}}(z^{\prime})$ is the critical density for spherical collapse at redshift $z^{\prime}$ extrapolated to the present using the linear theory. The term $\delta_{0}(\delta,z)$ corresponds to the initial density necessary for a region to reach $\delta$ at redshift $z$. Therefore, within each cell of volume $V_{\rm cell}$, we generate a random sample of low-mass halos within a specific mass bin, ensuring $n_{\rm cell}(m)$ follows a Poisson distribution with mean $n(m)$. Similar to resolved halos, these randomly generated halos are then converted into HI, and subsequently, into brightness temperature.
@@
where $f_{\mathrm{H}}=0.76$ is the hydrogen fraction.
After generating these observed brightness temperature data, we further perform data augmentation by rotating the datacube around a fixed line-of-sight direction as the rotation axis to generate additional samples. This process results in a total of $150$ samples of 21cm brightness temperature. The dataset is then split into three subsets: $80\%$ is used for training the U-Net models, $10\%$ is reserved for validation to fine-tune hyperparameters, and the remaining $10\%$ is used for testing to evaluate the model’s generalization and performance.
@@ 3.2 U-Net Architecture and Training
As shown in Figure 1, our U-Net model consists of five layers each of down-sampling and up-sampling process. In the down-sampling phase, the encoder incorporates convolution and pooling layers to gradually reduce the high spatial resolution of the simulated data. To process large volumetric data, we use the successive double convolutional layers of $3\times 3\times 3$ followed by a max pooling layer with $2\times 2\times 2$ kernel size to efficiently capture the three dimensional spatial features. In the up-sampling phase, transposed convolution layers of $2\times 2\times 2$ are utilized to restore spatial resolutions. Additionally, skip connections create direct links between down-sampling and up-sampling layers with matching corresponding resolutions, allowing easier flow of features and gradients during backpropagation, while enhancing the preservation of complex details and helping to alleviate computational overhead. The ReLU activation function is applied after each convolution layer, allowing the network to capture and retain effective feature information from deeper levels.
Figure 3: Comparison of the two-dimensional power spectrum $P(k_{\perp},k_{\parallel})$ of 21cm brightness temperature across various fields and resolutions. The first row displays the power spectrum of the true brightness temperature. The second row presents the observed map, with modes contaminated by foregrounds removed, specifically those with $k_{\parallel}<0.1~{}{\rm h/Mpc}$ and within the wedge. The last row displays the power restored by the U-Net model. Additionally, different columns correspond to power spectra calculated from maps at varying resolutions, starting from $64^{3}$ up to $320^{3}$. As detailed in the text, the AI model initially trained at a fixed grid size of $128^{3}$, has been effectively applied to varying resolutions.
@@
The training of neural networks begins with forward propagation, during which the network processes input data to extract features and produce predictions. Therefore the quality of this input data significantly impacts the network’s training efficacy and the accuracy of its predictions. Training neural networks requires considerable computational resources to update the weights and biases during training. Higher-resolution input data allows the model to capture finer details of LSS, which encode valuable information about non-linear mode coupling. However, this improvement comes at the cost of increased computational demands, as higher resolution results in larger input volumes and a corresponding increase in the number of hyperparameters that must be processed.
@@
To optimize model performance while managing computational constraints, we divided the extensive dataset into smaller batches and experimented with different resolutions to evaluate their effects on model efficacy. Our goal is to achieve an optimal balance between computational limitations and the maximization of usable data. From the previously described simulated 21cm brightness temperature data, we randomly selected $80$ percent as the training set, in which foreground contaminated modes have been excluded. Specifically we removed modes where $k_{\parallel}<0.3k_{\perp}$ and $k_{\parallel}<0.1{\rm Mpc/h}$ in Fourier space. After some testings, we opted to utilize these datacubes with a grid size of $128^{3}$. As discussed later, our AI model, trained with this lower resolution data, sufficiently captures the mode coupling features, and could be applied to predict higher resolution data and produce high-quality modes restoration.
@@
where $N$ represents the number of training samples in a single batch. This loss function quantifies the deviation of the model’s predictions from the true values for each batch. With above loss function, we employ the Adam optimizer [69] to adjust the model weights and biases by minimizing the squared errors during the training loop, with a learning rate of $10^{-3}$. The backpropagation process is initiated to fine-tune the network parameters, allowing the generated results to progressively converge with the true values. This iterative adjustment continues until the network achieves convergence.

### v2

@@ Abstract
In 21cm intensity mapping of the large-scale structure (LSS), regions in Fourier space could be compromised by foreground contamination. In interferometric observations, this contamination, known as the foreground wedge, is exacerbated by the chromatic response of antennas, leading to substantial data loss. Meanwhile, the baryonic acoustic oscillation (BAO) reconstruction, which operates in configuration space to "linearize" the BAO signature, offers improved constraints on the sound horizon scale. However, missing modes within these contaminated regions can negatively impact the BAO reconstruction algorithm. To address this challenge, we employ the deep learning model U-Net to recover the lost modes before applying the BAO reconstruction algorithm. Despite hardware limitations, such as GPU memory, our results demonstrate that the AI-restored 21cm temperature map achieves a high correlation with the original signal, with a correlation ratio of approximately $0.9$ at $k\sim 1{\rm h/Mpc}$. Furthermore, subsequent BAO reconstruction indicates that the AI restoration has minimal impact on the performance of the ‘linearized’ BAO signal, proving the effectiveness of the machine learning approach to mitigate the impact of foreground contamination. Interestingly, we demonstrate that the AI model trained on coarser fields can be effectively applied to finer fields, achieving even higher correlation ratio. This success is likely attributable to the scale-invariance properties of non-linear mode coupling in large-scale structure and the hierarchical structure of the U-Net architecture.
@@ 1 Introduction
This missing information can significantly impact cosmological measurements. For one, the loss of Fourier modes reduces the statistical constraining power.For baryonic acoustic oscillations (BAO), this loss also affects data post-processing. Most current BAO measurements in galaxy surveys utilize a technique known as BAO reconstruction [35, 36, 37, 38, 39, 40, 41, 42], which aims to mitigate the non-linear degradation of the BAO signal. Despite the technical differences among various reconstruction algorithms, these methods generally seek to solve the continuity equation to ‘linearize’ the BAO signature. However, contamination of a large fraction of Fourier modes can significantly impact the reconstruction outcome. [43] investigated this effect using Fisher information at redshifts $z=1-2$ and found that it could increase errors by a factor of $3-4$ for the angular diameter distance and by $\sim 1.5$ for $H(z)$. One potential solution is to use a low-density galaxy sample to compensate for the affected modes [44]. Alternatively, information loss can be mitigated by leveraging nonlinear mode coupling. For example, the tidal reconstruction method [45] employs a quadratic estimator derived from the tidal field to recover large-scale modes from small-scale perturbations. However, as a perturbative approach, its effectiveness is primarily limited to large scales.
With the widespread application of artificial intelligence (AI), machine learning (ML) has become a powerful tool for processing large-scale structure data and analyzing cosmic datasets [46, 47, 48, 49, 50, 51, 52, 53, 54]. It has also proven to be highly effective in addressing high-dimensional astronomical problems, such as reconstructing BAO signals from dark matter density fields [55], isolating the EoR signals with complex frequency-dependent beam effects [56], and identifying ionization regions in images with instrument noise [57, 58]. Therefore, in this study, we aims to apply the deep learning technique, particularly the U-Net model, to recover the large-scale structure modes in Fourier regions affected by foreground contamination. Additionally, we will examine the impact of the AI-restored data on BAO reconstruction. The structure of this paper is as follows: Section 2 provides a brief overview of foreground contamination and its impact on BAO reconstruction. In section 3, we introduce our main methodology, including details of the simulation data for 21cm intensity mapping, the U-Net architecture, and the BAO reconstruction process. The results are presented in Section 4, and we conclude with a discussion in Section 5.
@@ 2 Foreground Contamination and the Impact on BAO Reconstruction
In both single-dish and interferometric observations, the spectral smoothness of the foreground indicates that the LSS signal may be compromised within the Fourier region where the wavenumbers along the line-of-sight direction are small, i.e. $k_{\parallel}<k_{\rm int}$ where $k_{\parallel}$ represent the line-of-sight component of the wave vector $\mathbf{k}$. This is commonly referred to as the intrinsic foreground. The exact value of $k_{\rm int}$ depends on the specific foreground removal techniques employed and specific instrumental systematics; for our analysis, we have adopted a reasonable value of $k_{\rm int}=0.1~{}{\rm h/Mpc}$. Furthermore, the chromatic response of the interferometer complicates the situation [59, 60, 61] by introducing additional coupling between spectral and spatial information. This coupling can lead to foreground leakage from low $k_{\parallel}$ to high $k_{\perp}$, with the effect being particularly severe at high $k_{\perp}$, where $k_{\perp}$ is the transverse components of $k$.
@@
in the two-dimensional $k_{\parallel}-k_{\perp}$ plane. In this paper, we consider $\theta_{\rm FOV}=\pi/2$, at the redshift of $z=1$, and the cosmological parameters as follows $\Omega_{m}=0.31,\Omega_{b}=0.048,\Omega_{\Lambda}=0.69,\sigma_{8}=0.82,h=0.68$. The combination of intrinsic foreground and the mode mixing wedge results in significant information loss, which, as will be discussed, substantially impacts BAO reconstruction.
Despite its robustness and enormous success in constraining cosmological parameters in LSS surveys, the baryonic acoustic oscillation is still subject to various systematics. One of the most well-understood effects is the broadening of the BAO peak caused by the non-linear evolution of the bulk velocity of LSS [62]. In Fourier space, this manifests as a dampening of the BAO wiggles, which can be modeled by a simple exponential function. Compared to the linear BAO peak, this effect reduces the accuracy of measurement in a given LSS survey when extracting the sound horizon scale. Moreover, the same non-linear process also leads to the emergence of non-Gaussianity, which results in additional information leakage into higher-order statistics and further decreases the overall constraining power of the measurements. Various techniques have been developed to reverse such process. For example, [35] proposed a reconstruction algorithm based on the linear Zel’dovich approximation which can partially reverse nonlinear degradation of BAO signal. This technique has been widely used in recent galaxy surveys [63, 64, 65, 66]. Additionally, [40] extended this method to pixelated maps for applications in 21cm intensity mapping.
@@
where $\mathbf{q}$ and $\mathbf{x}$ are Lagrangian and Eulerian coordinates of particles respectively, $\delta^{K}_{ij}$ is the Kronecker delta, the displacement potential and vector are defined with $\mathbf{\Psi}(\mathbf{q})=\mathbf{x}-\mathbf{q}=\nabla\phi(\mathbf{q})$. By solving equation (2.4), the reconstruction algorithm reverses such a non-linear transformation and produces a reconstructed $\phi_{\rm rec}$ field on a grid that is close to the Lagrangian coordinates. In the following, we will denote the Laplacian of the reconstructed field $\delta_{\rm rec}=\nabla^{2}\phi_{\rm rec}$. Consequently, the BAO signature of the reconstructed field closely resembles that of the linear theory. Since these algorithms solve for the displacement field in the configuration space, in 21cm intensity mapping experiments, the missing modes due to the foreground contamination will inevitably impact the performance of the reconstruction.
@@ 3.1 Data Preparation
At low redshifts, neutral hydrogen primarily resides within galaxies. In this subsection, we describe the generation of training data for our AI model. Given the computational cost of producing large simulation datasets, we adopt several simplifications. First, We use the COLA (COmoving Lagrangian Acceleration) simulations code  [^1] https://bitbucket.org/tassev/colacode/src/hg/ [67], which integrate second-order Lagrangian Perturbation Theory (2LPT) with N-body algorithms, to rapidly simulate multiple realizations of dark matter distributions. Second, to assign HI mass $M_{\mathrm{HI}}$, we follow the fitting recipe from [16] to populate dark matter halos with neutral hydrogen. Due to the limited mass resolution, we also introduce a subgrid population to account for halos below the simulation’s resolution threshold.
For the dark matter simulation, we conducted 37 realizations using the same set of cosmological parameters described previously but with varying initial seeds. Each simulation contains $1024^{3}$ particles within a cubic box measuring $800{\rm h^{-1}Mpc}$. The initial conditions for these simulations were set at a redshift of $z_{\rm init}=29$. We then generate the corresponding distributions of neutral hydrogen for each realization. Given the relatively low mass resolution of our dark matter simulations, the resolved halo masses are significantly larger than the minimum halo mass that can physically host HI. Consequently, our HI map consists of two distinct components: resolved halos and a subgrid contribution. For resolved halos, we employ the Robust Overdensity Calculation using K-Space Topologically Adaptive Refinement (ROCKSTAR) halo finder [68] to identify halo catalogs. This includes determining their locations, masses, and peculiar velocities. Furthermore, we account for the redshift space distortions (RSD) by adjusting the real space coordinates of halos according to their peculiar velocities $v_{r}=\textbf{v}\cdot\hat{\textbf{r}}$, along the line-of-sight (LoS) direction, which we designate as the $z$ axis. The redshift space position s is then defined by
@@
For subgrid contribution, we assume that these low-mass halos trace the dark matter field. Therefore, with the position and the peculiar velocity of the dark matter particles, we construct the redshift space matter distribution using Equation (3.1). In each cell with volume $V_{\rm cell}$ and mass $M_{\rm cell}$ at redshift $z$, we assume that the number of halos with mass $m$ follows a Poisson distribution, with the average number density determined by the halo mass function. For halos that are viralized at redshift $z^{\prime}$, the average number density can be calculated by [69]
@@
$\nu_{10}=\frac{\left[\delta_{sc}(z^{\prime})-\delta_{0}(\delta,z)\right]^{2}}{\sigma^{2}(m)-\sigma^{2}(M_{\mathrm{cell}})}.$ (3.3)
Here $\rho_{\mathrm{crit}}$ is the critical density, $\sigma(m)$ represents the variance of density fluctuations on mass $m$ and $\delta_{\text{sc}}(z^{\prime})$ is the critical density for spherical collapse at redshift $z^{\prime}$ extrapolated to the present using the linear theory. The term $\delta_{0}(\delta,z)$ corresponds to the initial density necessary for a region to reach $\delta$ at redshift $z$. Therefore, within each cell of volume $V_{\mathrm{cell}}$, we generate a random sample of low-mass halos within a specific mass bin, ensuring that $n_{\mathrm{cell}}(m)$ follows a Poisson distribution with mean $n(m)$. Similar to resolved halos, these randomly generated halos are then converted into HI, and subsequently, into brightness temperature.
@@
where $f_{\mathrm{H}}=0.76$ is the hydrogen fraction. After generating these observed brightness temperature data, we further perform data augmentation by rotating the datacube around a fixed line-of-sight direction as the rotation axis to generate additional samples. This process results in a total of $37\times 4=148$ samples of 21cm brightness temperature. The dataset is then split into three subsets: $80\%$ is used for training the U-Net models, $10\%$ is reserved for validation to fine-tune hyperparameters, and the remaining $10\%$ is used for testing to evaluate the model’s generalization and performance.
However, our HI modeling has clear limitations. The empirical HI assignment does not fully capture its true clustering properties, which may impact our AI model by altering the underlying mode coupling. Additionally, the low mass resolution of our dark matter simulation requires subgrid modeling of HI content, introducing biases and failing to preserve the spatial correlations present in a realistic HI distribution. Furthermore, [16] showed that while the bulk velocity of HI closely follows that of its host halo in smaller halos, it deviates significantly for larger halos. Moreover, HI velocity dispersion is generally lower than that of dark matter, particularly in small halos. As a result, the redshift-space distortions in both resolved and unresolved halos in our model may be biased. Despite these drawbacks, this method remains a practical compromise for generating the necessary training data. In future work, we will focus on higher-resolution simulations and explore alternative HI modeling and velocity treatments to improve accuracy.
@@ 3.2 U-Net Architecture and Training
As shown in Figure 1, our U-Net model consists of five layers each of the down-sampling and up-sampling process. In the down-sampling phase, the encoder incorporates convolution and pooling layers to gradually reduce the high spatial resolution of the simulated data. To process large volumetric data, we use successive double convolutional layers of $3\times 3\times 3$ followed by a max pooling layer with $2\times 2\times 2$ kernel size to efficiently capture three-dimensional spatial features. In the up-sampling phase, transposed convolution layers of $2\times 2\times 2$ are utilized to restore spatial resolutions. Additionally, skip connections create direct links between down-sampling and up-sampling layers with matching corresponding resolutions, allowing easier flow of features and gradients during backpropagation, while enhancing the preservation of complex details and helping to alleviate computational overhead. The ReLU activation function is applied after each convolution layer, allowing the network to capture and retain effective feature information from deeper levels.
@@
The training of neural networks begins with forward propagation, during which the network processes input data to extract features and produce predictions. Therefore, the quality of this input data significantly impacts both the network’s training efficacy and the accuracy of its predictions. Training neural networks requires considerable computational resources to update the weights and biases during training. Higher-resolution input data allows the model to capture finer details of LSS, which encode valuable information about non-linear mode coupling. However, this improvement comes at the cost of increased computational demands, as higher resolution results in larger input volumes and a corresponding increase in the number of hyperparameters that must be processed.
@@
To optimize model performance while managing computational constraints, we divided the extensive dataset into smaller batches and experimented with different resolutions to evaluate their effects on model efficacy. Our goal is to achieve an optimal balance between computational limitations and the maximization of usable data. From the previously described simulated 21cm brightness temperature data, we randomly selected $80$ percent as the training set, in which foreground contaminated modes have been excluded. After some testing, we adopted a two-step training procedure to maximize performance within our GPU constraints, using the same U-Net architecture (Figure 1). In the first stage, we trained the AI model using simulated data cubes with a grid size of $128^{3}$ (stage-one model). This initial model already performed reasonably well, effectively capturing mode-coupling features and demonstrating strong capability in restoring high-quality modes, even when applied to higher-resolution data. However, it exhibited noticeable amplitude deviations in the deeply nonlinear regime for higher resolution maps. To further improve performance, we retrained this model using a smaller dataset of only $13$ simulated fields at a higher resolution of $256^{3}$ (stage-two model), again limited by available GPU resources.
@@
where $N$ represents the number of training samples in a single batch. This loss function quantifies the deviation of the model’s predictions from the true values for each batch. With the above loss function, we employ the Adam optimizer [72] to adjust the model weights and biases by minimizing the squared errors during the training loop, with a learning rate of $10^{-3}$. The backpropagation process is initiated to fine-tune the network parameters, allowing the generated results to progressively converge with the true values. This iterative adjustment continues until the network achieves convergence.